OpenStudy (anonymous):

(Will give cookie and medal :D) Solve for x: x - 6 ---- = 5 4

OpenStudy (anonymous):

Sorry, it's a little hard to get the question is it \[\frac{x-6}{4} = 5\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok great. What do you think the first step would be? We are trying to solve for x and we want to get rid of the numbers on the bottom of the fraction.

OpenStudy (anonymous):

so we'd get rid of the fraction first?

OpenStudy (anonymous):

Yep, we need to multiply both sides by something to get rid of the fraction. What do you think we should multiply by?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

BOOM! that's it! Multiplying both sides by 4 gives us \[\ 4\times \frac{x-6}{4} = 5 \times 4\]Can you simplify that?

OpenStudy (anonymous):

lol... uhh so multiplying 6 and 4 first?

OpenStudy (anonymous):

Well we need to look at the right hand side like this \[4 \times (\frac{x-6}{4}) = x-6\] When we multiply that out we can simply cancel out. Does that make sense?

OpenStudy (anonymous):

im a bit confused. I thought we just multiply all the numbers in the equation by 4 making 6 24 and so on

OpenStudy (anonymous):

I'd just convert 5 to a fraction with a denominator of 4, cancel the denominators and solve for x.

OpenStudy (anonymous):

as in 5/4 x 4/1?

OpenStudy (anonymous):

Yep that works too, however it's a little faster this way (and easier in my opinion. However we can do that. \[4 \times (\frac{x-6}{4}) = \frac{4x-24}{4} = \frac{4x}{4}-\frac{24}{4}\] \[\frac{4x}{4}-\frac{24}{4} = x-6\]So hopefully you can see that we reach the same conclusion but it takes a lot longer.

OpenStudy (anonymous):

no. As in: (x-6)/4 = 20/4 cancelling the denominators leaves x-6 = 20

OpenStudy (anonymous):

ohhh... so x would equal 26?

OpenStudy (anonymous):

Yup... :-)

OpenStudy (anonymous):

oh okay thanks you guys!!! *gives each a cookie*

OpenStudy (anonymous):

Danke!! :-)

OpenStudy (anonymous):

Cheers!

OpenStudy (anonymous):

69

OpenStudy (anonymous):

soryy